OpenStudy (anonymous):

Solve x2 = 14x – 49. I'm really confused.... It has to be factored and I can't figure out how... :( Please help!

OpenStudy (phi):

can you put it into "standard form" ?

OpenStudy (anonymous):

x2 - 14x - 49 = 0 or -x2 + 14x - 49 = 0

OpenStudy (anonymous):

* x2 - 14x + 49 = 0

OpenStudy (anonymous):

That looks good. Could you use the Quadratic formula from here?

OpenStudy (phi):

yes. now list all the ways to multiply 2 numbers that give you 49 1,49 7,7 those are your only choices which pair can be added to give you 14 ?

OpenStudy (anonymous):

7 and 7 but they don't multiply to -49.... You would have to multiply -7 and 7 and that would give you a sum of 0....

OpenStudy (phi):

x^2 - 14x + 49 = 0 the + on the 49 means the factors will have the same sign the - on the 14 means the larger factor is - that means both are -

OpenStudy (anonymous):

okay

OpenStudy (phi):

I am not sure why you say ***7 and 7 but they don't multiply to -49 unless you were not looking at the equation in standard form, where you have a +49

OpenStudy (anonymous):

I was looking at the other version of standard form... You can either move the x2 or the other part and I looked at the wrong one.... So x=7?

OpenStudy (phi):

yes, x=7 as a check: x^2 = 14x – 49 ? 7*7 = 14*7 - 49 ? 49 = 98 -49 ? 49 = 49 yes. it worked